5/12 , 3/20 find the lcd then c onvert each rational expression into an equivelant rational expression with the lcd as the denominator
Found 2 solutions by stanbon, edjones:
Answer by stanbon(75887)

5/12 , 3/20 find the lcd then c onvert each rational expression into an equivelant rational expression with the lcd as the denominator
----------------
The lcd of 12 and 20 is 60
5/12 = (5*5)/(12*5) = 25/60 = 25/lcd
---------
3/20 = (3*3)/(20*3) = 9/60 = 9/lcd
===================
Cheers,
Stan H.

Answer by edjones(8007)

12,24,36,48,60
20,40,60
25/60,9/60
the lcd or lcm (least common multiple)is the lowest number that both denominators go into evenly (ie. without leaving a remainder).
To get the correct numerator divide the numerator into the lcd and multiply the result times the old numerator to get the new numerator. that gives the new number the same RATIO as the old number.
Rational comes from the word ratio.
Now you can add or subtract the numbers easily.(25/60+9/60=34/60=17/30)(25/60-9/60=16/60=4/15)
Ed